### v1

# On Signal Peak Power Constraint of Over-the-Air Federated Learning

Abstract
Federated learning has been considered a promising privacy preserving distributed edge learning framework. Over-the-air computation technique leveraging analog transmission enables the aggregation of local updates directly over-the-air by exploiting the superposition properties of wireless multiple-access channel, thereby drastically reducing the communication bottleneck issues of federated learning compared with digital transmission schemes. This work points out that existing over-the-air computation-federated learning overlooks a key practical constraint, the instantaneous peak-power constraints imposed by the non-linearity of radio-frequency power amplifiers. We present and analyze the effect of the classic method to deal with this issue, amplitude clipping combined with filtering. We investigate the effect of instantaneous peak-power constraints in over-the-air computation-federated learning for both single-carrier and multi-carrier orthogonal frequency-division multiplexing systems. We highlight the specificity of over-the-air computation-federated learning: the samples depend on the gradient value distribution, leading to a higher peak-to-average power ratio than that observed for uniformly distributed signals. Simulation results demonstrate that, in practical settings, the instantaneous transmit power regularly exceeds the power-amplifier limit; however, by applying clipping and filtering, the FL performance can be degraded. The degradation becomes pronounced especially in multi-carrier OFDM systems due to the in-band distortions caused by clipping and filtering.
Index Terms:
federated learning, over-the-air computation, peak-to-average, peak power constraint.
I Introduction
An unprecedented volume of data is now generated at the wireless edge. Leveraging these data to train machine learning models at a centralized server is challenged by privacy concerns and limited uplink bandwidth requirements. Federated learning addresses both challenges by keeping raw data on devices and only exchanging model updates [1]. Federated learning consists of a collaborative learning framework, in which repeatedly, each edge device trains on its own data for certain iterations and then uploads the model or the gradients to a central server for aggregation. However, the digital transmission of large-scale model updates still creates a communication bottleneck, especially as the number of participants scales [2]. As more and more end devices are wirelessly connected, considering federated learning within wireless networks is essential [3].
Over-the-air computation technique proposed in [4, 5] is investigated as a means to counter the communication bottleneck issue of federated learning [6, 7, 8]. By exploiting the waveform superposition of the wireless multiple-access channel, a server can directly aggregate local updates “over the air,” enabling scalability with the number of devices to participate in each communication round. One of the central design challenges in over-the-air computation-federated learning is power control for signal power or amplitude alignment. The ideal aim is to obtain equally average received power from user equipments at the base station so that a summation operation can be performed. Under average power constraints, channel inversion and threshold-based policies have been shown to minimize aggregation mean-squared error: user equipments with weak channels transmit at their maximum power budget, while user equipments above a threshold invert their channels [9, 10, 11].
However, the above designs ignore the instantaneous peak-power limitations due to the non-linearity of power amplifiers from a certain power level. Operating in the nonlinear region causes signal distortion and receiver misalignment. This is a common issue for orthogonal frequency-division multiplexing, whose peak-to-average power ratio is known to be high. While recent works have advanced over-the-air computation in orthogonal frequency-division multiplexing systems [12, 13, 14], they typically optimize mean-squared error under average-power budgets and imperfect channel state information but do not quantify or enforce instantaneous peak-power constraints.
This paper studies over-the-air computation-federated learning under both average and instantaneous peak-power constraints for single-carrier transmission and multi-carrier orthogonal frequency-division multiplexing systems. We characterize the induced aggregation error when peaks are constrained, and evaluate the effect of practical mitigation mechanisms.
The main contributions are summarized as follows:
  • We highlight the practical limitations of peak power constraint in a practical system due to the non-linear region of power amplifiers.
  • We present and evaluate the most common method to handle the peak power constraints: amplitude clipping that repeatedly applies clipping, i.e., clip the time-domain peak power value that is set by the nonlinearity of the power amplifier, and filtering the out-of-band distortion. Implementation details for both multiple-access schemes are provided.
  • Empirical evaluation and design insights. Evaluating under LeNet trained on CIFAR-10, we quantify peak-to-average power ratio, aggregation mean-squared error, and learning accuracy. We observe substantially high peak-to-average power ratio of a over-the-air computation-federated learning signal, either in single-carrier or multi-carrier OFDM. Enforcing peak constraints via clipping yields accuracy degradation and may cause model divergence.
The remaining sections are organized as follows: Section II introduces the FL-OTA system models for single-carrier and multi-carrier OFDM and the associated average power allocation based on MSE minimization. Section III formalizes the instantaneous peak-power constraint and presents the clipping mechanisms. Section IV presents the simulation settings and results.
II System Model
II-A Federated Learning
Consider a single-antenna base station serving a set $\mathcal{K}\triangleq\{1,\ldots,K\}$ of $K$ user equipments, where federated learning is performed. A supervised learning task is considered for FL. User equipment $k$ holds a local dataset $\mathcal{D}_{k}$ of size $D_{k}$. The local empirical loss function can be written as
$F_{k}(\bm{{w}})=\frac{1}{D_{k}}\sum_{x\in\mathcal{D}_{k}}\ell(\bm{{w}};x),$ (1)
where $\ell(\bm{{w}};x)$ denotes the loss of the prediction of the sample $x\in\mathcal{D}_{k}$ evaluated with model weights $\bm{{w}}\in\mathbb{R}^{N}$. Federated learning has the objective of minimizing the empirical global loss function:
$\min_{\bm{{w}}\in\mathbb{R}^{N}}\sum_{k\in\mathcal{K}}D_{k}F_{k}(\bm{{w}}).$ (2)
Equal size of local dataset is assumed, i.e. $D_{k}=D$ for any $k\in\mathcal{K}$. Federated learning consists of multiple communication rounds where $t$-th communication round is described as follows:
  1. 1.
    Base station broadcasts the global model from previous communication round $\bm{{w}}^{(t-1)}$ to all user equipments. This step is assumed without issue since base station has sufficient transmit power.
  2. 2.
    Each user equipment $k\in\mathcal{K}$ performs a mini-batch stochastic gradient descent step based on $w^{(t-1)}$ to minimize the local empirical loss $F_{k}$ over its local dataset $\mathcal{D}_{k}$. The resulting gradient vector is denoted as $\overline{\bm{{g}}}_{k}^{(t)}$.
  3. 3.
    When all user equipments are ready for uplink transmission, a pilot signal is sent orthogonally from each user equipment to base station (assuming sufficient resources for short pilot signals). We assume perfect channel state information at the base station as commonly assumed in the literature [8, 6]. With the channel state of each user equipment, base station determines the transmit power that each user equipment should use for over-the-air computation, and sends the power information to each user equipment.
  4. 4.
    All user equipments apply the same normalization that transforms $\overline{\bm{{g}}}_{k}^{(t)}$ to $\bm{{g}}_{k}^{(t)}$ so that $\operatorname{\mathbb{E}}[\bm{{g}}_{k}^{(t)}]=0$ and $\operatorname{\mathbb{E}}[\|\bm{{g}}_{k}^{(t)}\|^{2}_{2}]=N$ [8, 15]:
    $\bm{{g}}_{k}^{(t)}=\frac{\overline{\bm{{g}}}_{k}^{(t)}-\mu}{\Gamma},\vskip-2.84544pt$ (3)
    where the normalizing mean $\mu^{(t)}=\frac{1}{|\mathcal{S}_{c}|}\sum_{k\in\mathcal{S}_{c}}\mu_{k}^{(t)}$ and standard deviation $\Gamma^{(t)}=\frac{1}{|\mathcal{S}_{c}|}\sum_{k\in\mathcal{S}_{c}}\Gamma_{k}^{(t)}$ are obtained by averaging the statistics of each participating clients:
    $\mu_{k}^{(t)}=\frac{1}{m}\sum_{i=1}^{d}\overline{g}_{k}^{(t)}[i],\quad\Gamma_{k}^{(t)}=\sqrt{\frac{1}{m}\sum_{i=1}^{d}\overline{g}_{k}^{(t)2}[i]-\mu_{k}^{(t)2}},$ (4)
    where $\overline{\bm{{g}}}_{k}^{(t)}=(\overline{g}_{k}^{(t)}[i])_{i=1,\ldots,N}$ the gradient vector for end-device $k$. Note that $\mu^{(t)}$ and $\Gamma^{(t)}$ have to be uniform across all end-devices due to the nature of AirComp and must be estimated before the data communication.
  5. 5.
    Each then sends the normalized gradient $\bm{{g}}_{k}^{(t)}$ via analog signal with the power control information given by base station. The results via over-the-air computation at the base station gives the approximation of the gradient average $\tilde{\bm{{g}}}^{(t)}$:
    $\tilde{\bm{{g}}}^{(t)}\approx\overline{\bm{{g}}}^{(t)}\triangleq\frac{1}{K}\sum_{k\in\mathcal{K}}\overline{\bm{{g}}}_{k}^{(t)},\vskip-2.84544pt$ (5)
    (details will be shown in later section), and update the new global model $\bm{{w}}^{(t)}=\bm{{w}}^{(t-1)}-\eta\overline{\bm{{g}}}^{(t)}$ with $\eta>0$ the learning rate and $\overline{\bm{{g}}}^{(t)}$ the de-normalized gradients.
In the following sections, we focus on step 5 that consists of the gradient transmission via over-the-air computation techniques in single-carrier and multi-carrier OFDM-based systems. For ease of notation, the index of communication round $t$ is removed from all notations, since the principle is identical given any channel states and gradient vectors.
II-B Single-Carrier over-the-air computation
II-B1 Transmission of gradient
In the case of single-carrier transmission, each communication round is divided into $N$ time slots, where $N$ is the size of the trainable parameters of the neural network. In time slot $n=1,\ldots,N$, all devices transmit their $n$-th normalized gradient value $\bm{{g}}_{k}[n]$. Through the use of over-the-air computation, the superimposed characteristics of the wireless medium are exploited to calculate the target function (5). We assume block-fading, so that the channel characteristics stay constant over the transmission of one gradient vector as assumed in most works [6, 8]. The received signal at the centralized server in a given communication round can be written as
$\bm{{y}}=\sum_{k\in\mathcal{K}}h_{k}\sqrt{p_{k}}\bm{{g}}_{k}+\bm{{n}},$ (6)
with $h_{k}>0$ the channel gain between BS and user equipment $k$, $p_{k}$ the “average” transmit power of UE $k$, $\bm{{n}}\in\mathbb{R}^{N}$ the additive zero-mean Gaussian noise of variance $\sigma^{2}$.
Base station rescales the received superposed signal by a factor $\sqrt{\alpha}>0$ and dividing by the number of UEs $K$, and then de-normalize the gradient to obtain the $\tilde{\bm{{g}}}\in\mathbb{R}^{N}$, the estimation of (5):
$\tilde{\bm{{g}}}=\frac{\Gamma\sqrt{\alpha}\bm{{y}}}{K}+\mu.\vskip-2.84544pt$ (7)
II-B2 Power Control
Over-the-air computation inevitably faces noises and errors. The power $p_{k}$ and scaling factor $\alpha$ are controlled to minimize the mean-squared error between the AirComp gradient $\tilde{\bm{{g}}}$ (7) and the true gradient $\overline{g}$ (5) as
$\mathrm{MSE}$ $=\operatorname{\mathbb{E}}[\|\tilde{\bm{{g}}}-\overline{\bm{{g}}}\|_{2}^{2}]$ (8)
[…]
$=\frac{\Gamma^{2}N}{K^{2}}\Big(\sum_{k\in\mathcal{K}}(h_{k}\sqrt{\alpha p_{k}}-1)^{2}+\alpha\sigma^{2}\Big).$
The equation can be derived by independence of zero-mean Gaussian noise with $\bm{{g}}_{k}$ and that $\operatorname{\mathbb{E}}[\|\bm{{g}}_{k}\|_{2}^{2}]=N$ and $\operatorname{\mathbb{E}}[\bm{{g}}_{k}]=0$. A similar expression with a factor $N+1$ can also be obtained without such assumptions on the gradient vector, which can be obtained via the Cauchy-Schwarz inequality as shown in [16].
To optimize the MSE, we optimize the term in the outer bracket subject to the average power constraint $P_{\mathrm{avg},\max}$. The resulting optimization problem is given by
$\min_{\alpha>0,\ \{p_{k}\}_{k\in\mathcal{K}}}$ $\sum_{k\in\mathcal{K}}(h_{k}\sqrt{\alpha p_{k}}-1)^{2}+\alpha\sigma^{2}$ (9)
$\mathrm{s.t.}$ $(\forall k\in\mathcal{K})\quad 0\leq p_{k}\leq P_{\mathrm{avg},\max}.$
The optimal solution to the optimization problem has been developed in [16]. Later, it was also extended to the case with non-zero-mean gradient, resulting in an additional composite misalignment error in the MSE expression in [9]. The optimal solution has a structure of a threshold-based scheme where devices with weaker channels transmit at their maximum average power $P_{\mathrm{avg},\max}$, while those with channels exceeding the threshold use the respective channel inverting power.
Proposition 1 ([16, 9]).
Assuming $h_{1}\geq h_{2}\geq\ldots\geq h_{K}$. There exists $k^{*}\in\mathcal{K}$ such that the optimal solution $(\alpha^{*},(p_{k}^{*})_{k\in\mathcal{K}})$ to the problem (9) satisfies that the corresponding denoising factor is given by
$\alpha^{*}=\frac{1}{P_{\mathrm{avg},\max}}\biggl(\frac{\sum\limits_{k\geq k^{*}}h_{k}}{\sum\limits_{k\geq k^{*}}h_{k}^{2}+\frac{\sigma^{2}}{P_{\mathrm{avg},\max}}}\biggr)^{2},$ (10)
and the optimal average transmit power is given by
[…]
II-C Multi-Carrier OFDM-based over-the-air computation
II-C1 Transmission of gradient
In multi-carrier OFDM systems, individual values of $\bm{{g}}_{k}$ are modulated on subchannels in the frequency domain by assigning one value of $\bm{{g}}_{k}$ to one subchannel. Since the number of subchannels $M$ within the system bandwidth $B$ is typically less than the size of model parameter $N$, transmission spans multiple OFDM symbols. Let the OFDM symbol index be $\ell\in\mathcal{S}\triangleq\{0,\ldots,L-1\}$ with $L=\lceil\frac{N}{M}\rceil$. For $\ell$-th symbol, the subcarrier $m\in\mathcal{M}\triangleq\{0,\ldots,M-1\}$ carries the gradient entry of index $n=\ell M+m$. Zero values are transmitted for $n\geq N$. The $i$-th transmitted OFDM-symbol discrete sample in the time domain of the $k$-th user equipment and $\ell$-th OFDM-symbol is given by its IDFT:
$(\forall i\in\mathcal{M})\quad s_{k,\ell}[i]=\sum_{m=0}^{M-1}g_{k}[\ell M+m]b_{k,\ell M+m}e^{j\frac{2\pi mi}{M}},\vskip-2.84544pt$ (12)
where $b_{k,\ell M+m}\in\mathbb{C}$ scales the transmitted power to mitigate the effect of the channel. The received frequency domain signal is given by the transmitted signal $s_{k,\ell}[i]$ convolved with the channel impulse response $h_{k,\ell}[i]$ with the length $C+1$. We suppose that each device transmits its gradient perfectly synchronized, so that the received signal at the sample $i$, such that $C\leq i\leq M-1$ becomes
$y[i]$ $=\sum_{k\in\mathcal{K}}s_{k,\ell}[i]\circledast h_{k,\ell}[i]+n[i],$ (13)
$\hskip-22.76228pt=\sum_{k\in\mathcal{K}}\!\Big(\!\sum_{m=0}^{M-1}\!g_{k}[\ell M\!+\!m]b_{k,\ell M+m}e^{j\frac{2\pi mi}{M}}\!\Big)\!\circledast\!h_{k,\ell}[i]\!+\!n[i],$
[…]
$Y_{\ell M+m}$ $=\mathrm{DFT}\{\{y[i]\}_{i\in\mathcal{M}}\}[m]$ (14)
$\hskip-25.6073pt=\frac{1}{M}\sum_{k\in\mathcal{K}}\!\sum_{i=0}^{M-1}\!\Big(\!\sum_{m^{\prime}=0}^{M-1}\!g_{k}[\ell M\!+\!m^{\prime}]b_{k,\ell M+m^{\prime}}e^{j\frac{2\pi m^{\prime}i}{M}}\!\circledast\!h_{k,\ell}[i]$
$\quad\quad+n[i]\Big)e^{-j\frac{2\pi mi}{M}}$
$\hskip-25.6073pt=\sum_{k=1}^{K}g_{k}[\ell M+m]b_{k,\ell M+m}H_{k,\ell M+m}+N[m],$
where $H_{k,\ell M+m}=|H_{k,\ell M+m}|e^{j\phi_{k,\ell M+m}}$ is the frequency response of $h_{k,\ell}[i]$ and $N[m]\sim\mathcal{C}\mathcal{N}(0,\sigma^{2})$ is the additive white Gaussian noise for each sub channel since the Fourier transform of a Gaussian distribution is also a Gaussian distribution with the same statistical properties. Due to the fact that each sub channel is associated with a different channel $H_{k,\ell M+m}$ in OFDM, the $n$-th gradient value can therefore be recovered by:
$\tilde{g}[n]$ $=\frac{\Gamma\sqrt{\alpha_{n}}Y_{n}}{K}+\mu,$ (15)
$\hskip-14.22636pt=\frac{\Gamma\sqrt{\alpha_{n}}}{K}\sum_{k\in\mathcal{K}}|H_{k,n}|\sqrt{p_{k,n}}g_{k}[n]+\frac{\Gamma\sqrt{\alpha_{n}}}{K}N[n]+\mu,$
where $b_{k,\ell M+m}$ in (14) is substituted as $b_{k,\ell M+m}=\sqrt{p_{k,\ell M+m}}e^{-j\phi_{k,\ell M+m}}$, with $p_{k,\ell M+m}$ the transmit power of UE $k$ at the subcarrier $m$ when transmitting the $\ell$-th symbol, and hence already includes the phase offset caused by the channel and $\alpha_{n}>0$ is the denoising factor for the corresponding subcarrier.
II-C2 Power Allocation
As for single-carrier, the power allocation is done by minimizing the MSE of each gradient value. The MSE of the gradient value $n$ has the following expression:
$\mathrm{MSE}_{n}$ $=\operatorname{\mathbb{E}}[\|\tilde{g}[n]-\overline{g}[n]\|^{2}]$ (16)
$=\frac{\Gamma^{2}}{K^{2}}\Big(\sum_{k\in\mathcal{K}}(|H_{k,n}|\sqrt{\alpha_{n}p_{k,n}}-1)^{2}+\alpha_{n}\sigma^{2}\Big),$
For any OFDM-symbol $\ell$, the optimal power allocation for gradient values of indices $n=\ell M+m$ with $m=0,\ldots,M-1$ can be given as:
$\min_{\begin{subarray}{c}\bm{{\alpha}}_{\ell}\geq 0\\ (\bm{{p}}_{\ell M+m})_{m\in\mathcal{M}}\end{subarray}}$ $\sum_{m=0}^{M-1}\mathrm{MSE}_{\ell M+m}(\alpha_{\ell M+m},\bm{p}_{\ell M+m})$ (17a)
s.t. $\hskip-22.76228pt(\forall k\in\mathcal{K})\quad\sum_{m=0}^{M-1}p_{k,\ell M+m}\leq P_{\mathrm{avg},\max},$ (17b)
$\hskip-22.76228pt(\forall k\in\mathcal{K})\ (\forall m\in\mathcal{M})\ p_{k,\ell M+m}\geq 0,$ (17c)
with $\bm{\alpha}_{\ell}=(\alpha_{\ell M+m})_{m\in\mathcal{M}}$ and $\bm{{p}}_{n}=(p_{k,n})_{k\in\mathcal{K}}$ for any $n=\ell M+m$. This problem has been analyzed in [13] that requires a subgradient solver and partly in [14] that considers an equal $\alpha$ for all subcarriers.
Due to the short time frame that the power needs to be allocated, only low-complexity solution is practically feasible. Since transmit power at the UE side is typically low, we pick up the classical empirical observations that the power allocation for uplink multi-carrier OFDM has a marginal effect on the performance [17, 18]. Assuming equal power allocated to each subcarrier for each UE, i.e., replacing the constraint (17b) by
$(\forall k\in\mathcal{K})\ (\forall m\in\mathcal{M})\ p_{k,\ell M+m}\leq\frac{P_{\mathrm{avg},\max}}{M}.\vskip-4.26773pt$ (18)
The problem can therefore be decoupled for each subcarrier $m$, and the optimal solution for each subcarrier $m$ given a symbol $\ell$, i.e., given $(H_{k,\ell M+m})_{k\in\mathcal{K}}$, $P_{\mathrm{avg},\max}/M$ as the maximum average transmit power, and $\sigma^{2}$ as the noise power level, can be calculated again with Proposition 1.
III Instantaneous peak power constraint
Figure 1: Typical input and output power characteristics curve for a power amplifier.
Practical radio-frequency power amplifiers operate linearly only up to certain input power [19, 20]. Given a certain average power as input $P_{\mathrm{avg}}^{(in)}$, the actual operating region would be between zero and the maximum peak instantaneous input power which can be obtained by multiplying $P_{\mathrm{avg}}^{(in)}$ with peak-to-average power ratio (addition in dB domain), as illustrated in Fig. 1. Operating in nonlinear region leads to signal distortions, resulting in strong in-band and out-of-band distortions (details see later), which is to be avoided. Note that all previously determined powers are actually output powers, but can be easily converted to the input power by a one-to-one relationship. We denote $p_{k}^{(in)}$ and $b_{k\ell M+m}^{(in)}$ respectively the corresponding input power of $p_{k}$ and $b_{k\ell M+m}$. We model this by an instantaneous amplitude limit $A_{\max}>0$ that corresponds equivalently to the input peak-power limit $P_{\mathrm{inst},\max}^{(in)}=A_{\max}^{2}$. The transmitted time-domain waveform $X(t)$ generated at a user equipment must satisfy
$|X(t)|^{2}\leq P_{\mathrm{inst},\max}^{(in)}=A_{\max}^{2}.\vskip-2.84544pt$ (19)
III-A Peak-to-Average Power Ratio
This limitation directly interacts with the peak-to-average power ratio of the transmitted waveform as illustrated in Fig. 1. A low PAPR can keep the operating region within the linear region.
Single-carrier transmission
With the gradient normalization in Section II, the average per-communication round transmit power equals the configured average power budget. The peaks for single carrier transmission are directly determined by the gradient values. The $k$-th user equipment transmits the real-valued sample $g_{k}[n]$ matching the amplitude of waveform. Its PAPR is
$\mathrm{PAPR}_{k}=\frac{\max_{n}|g_{k}[n]|^{2}}{\frac{1}{N}\sum_{n=1}^{N}|g_{k}[n]|^{2}}.\vskip-2.84544pt$ (20)
Multi-carrier OFDM transmission
The time-domain signal for multi-carrier OFDM can be written as
$s_{k,\ell}(t)=\sum_{m=0}^{M-1}g_{k}[\ell M+m]b_{k,\ell M+m}e^{j\frac{2\pi mt}{M}}.\vskip-4.26773pt$ (21)
The ideal PAPR is denoted as
$\mathrm{PAPR}_{k,\ell}^{\mathrm{ideal}}=\frac{\max_{t}|s_{k,\ell}(t)|^{2}}{\frac{1}{T}\int_{0}^{T}|s_{k,\ell}(t)|^{2}dt}.$ (22)
In practice, only digital discrete samples can be employed to estimate PAPR. It has been shown that samples sampled at Nyquist rate inaccurately captures PAPR.
Oversampling: To accurately capture PAPR in the time-domain OFDM signal, oversampling is essential. An oversampling factor $L_{\mathrm{os}}$ (typically $L_{\mathrm{os}}=4$ is sufficient [21]) is applied by zero-padding the frequency-domain signal before the IDFT. The oversampled time-domain signal is generated as:
$s_{k,\ell}^{(\mathrm{os})}[i]=\!\!\!\sum_{m=0}^{L_{\mathrm{os}}M-1}\!\!G_{k,\ell}[m]e^{j\frac{2\pi mi}{L_{\mathrm{os}}M}},\quad i\!=\!0,\ldots,L_{\mathrm{os}}M\!-\!1,\vskip-2.84544pt$ (23)
where the zero-padded frequency-domain vector is defined as:
$\hskip-7.11317ptG_{k,\ell}[m]\!=\!\begin{cases}g_{k}[\ell M\!+\!m]b_{k,\ell M+m}^{(in)},&0\!\leq\!m\!<\!M,\\ 0,&M\!\leq\!m\!<L_{\mathrm{os}}M.\end{cases}\vskip-2.84544pt$ (24)
The per-symbol PAPR can therefore be estimated as
$\mathrm{PAPR}_{k,\ell}=\frac{\max_{i}|s_{k,\ell}^{(\mathrm{os})}[i]|^{2}}{\frac{1}{L_{\mathrm{os}}M}\sum_{i=0}^{L_{\mathrm{os}}M-1}|s_{k,\ell}^{(\mathrm{os})}[i]|^{2}}.\vskip-4.26773pt$ (25)
III-B Clipping with Oversampling
The commonly used technique to deal with the high peak-to-average power ratio issue is applying the procedure of iterative clipping-and-filtering. Since clipping results in in-band and out-of-band distortions in multi-carrier OFDM [19, 20]. A low-pass filter is applied afterwards to limit out-of-band distortion, which may result in the rise of PAPR, necessitating repeated such procedures.
Amplitude Clipping
Let $\mathcal{C}_{A_{\max}}(x)\triangleq\min\{1,A_{\max}/|x|\}\,x$ denote the time-sample-wise amplitude clipper.
III-B1 Single-carrier
Each UE transmits
$x_{k}[n]\!=\!\mathcal{C}_{A_{\max}}\!\bigl(\!\sqrt{\!p_{k}^{\!(in)}}\!\!g_{k}[n]\!\bigr)\!=\!\min\Bigl\{1,\!\frac{A_{\max}}{\sqrt{\!p_{k}^{\!(in)}}\,\!\!\!g_{k}[n]}\!\Bigr\}\!\sqrt{p_{k}^{\!(in)}}\,\!\!\!g_{k}[n].$ (26)
In single-carrier, this is equivalent to considering the transmission of a clipped version of the gradient values, with nothing else altered by the clipping.
III-B2 Multi-Carrier OFDM
Amplitude Clipping
For each OFDM symbol $\ell$, the oversampled complex time-domain samples are clipped while preserving phase:
$x_{k,\ell}^{(\mathrm{os})}[i]=\mathcal{C}_{A_{\max}}\bigl(s_{k,\ell}^{(\mathrm{os})}[i]\bigr)=\min\Bigl\{1,\frac{A_{\max}}{|s_{k,\ell}^{(\mathrm{os})}[i]|}\Bigr\}s_{k,\ell}^{(\mathrm{os})}[i].$ (27)
Since the clipping operation is nonlinear, distortions are inevitable. The clipping distortion in the time domain can be written as: $d_{k,\ell}^{(\mathrm{os})}[i]=x_{k,\ell}^{(\mathrm{os})}[i]-s_{k,\ell}^{(\mathrm{os})}[i].$
Figure 2: Flow Diagram for iterative clipping-and-filtering in multi-carrier OFDM systems.
Frequency-Domain Distortion
Transforming the clipped signal back to the frequency domain via DFT yields:
$X_{k,\ell}[m]=\mathrm{DFT}\bigl\{x_{k,\ell}^{(\mathrm{os})}[i]\bigr\}[m]=G_{k,\ell}[m]+D_{k,\ell}[m],$ (28)
where $D_{k,\ell}[m]=\mathrm{DFT}\bigl\{d_{k,\ell}^{(\mathrm{os})}[i]\bigr\}[m]$ represents the clipping distortion in the frequency domain.
This distortion consists of two components:
  • In-band distortion: distortion on the considered subcarriers $m\in\mathcal{M}$, which degrades the aggregated gradients.
  • Out-of-band emissions: spectral regrowth on the zero-padded subcarriers $M\leq m<L_{\mathrm{os}}M$ due to clipping, which causes interference to adjacent channels.
Filtering
To suppress out-of-band emissions, a low-pass filteris applied. Define a rectangular filter:
$H_{\mathrm{filter}}[m]=\begin{cases}1,&0\leq m<M,\\ 0,&M\leq m<L_{\mathrm{os}}M.\end{cases}\vskip-2.84544pt$ (29)
The filtered frequency-domain signal is:
$X_{k,\ell}^{(\mathrm{filt})}[m]=X_{k,\ell}[m]\cdot H_{\mathrm{filter}}[m].\vskip-2.84544pt$ (30)
Converting back to the time domain yields the filtered signal:
$x_{k,\ell}^{(\mathrm{filt})}[i]=\mathrm{IDFT}\bigl\{X_{k,\ell}^{(\mathrm{filt})}[m]\bigr\}[i].\vskip-4.26773pt$ (31)
However, filtering may reintroduce peak regrowth, as it restores some of the constructive interference among subcarriers that clipping had disrupted [19].
Iterative Clipping and Filtering
To jointly control PAPR and out-of-band emissions, an iterative clipping-and-filtering procedure is commonly employed [22]. The algorithm proceeds as follows for each symbol $\ell$:
  1. 1.
    Initialize: $x_{k,\ell}^{(0)}[i]=s_{k,\ell}^{(\mathrm{os})}[i]$; $j=1$.
  2. 2.
    While $x_{k,\ell}^{(j,\mathrm{clip})}[i]$ has out-of-band distortion higher than a threshold:
    1. (a)
      Clip: $x_{k,\ell}^{(j,\mathrm{clip})}[i]=\mathcal{C}_{A_{\max}}\bigl(x_{k,\ell}^{(j-1)}[i]\bigr)$.
    2. (b)
      DFT: $X_{k,\ell}^{(j)}[m]=\mathrm{DFT}\bigl\{x_{k,\ell}^{(j,\mathrm{clip})}[i]\bigr\}[m]$.
    3. (c)
      Filter: $X_{k,\ell}^{(j,\mathrm{filt})}[m]=X_{k,\ell}^{(j)}[m]\cdot H_{\mathrm{filter}}[m]$.
    4. (d)
      IDFT: $x_{k,\ell}^{(j)}[i]=\mathrm{IDFT}\bigl\{X_{k,\ell}^{(j,\mathrm{filt})}[m]\bigr\}[i]$.
  3. 3.
    Output: $x_{k,\ell}^{(\mathrm{ICF})}[i]=x_{k,\ell}^{(J_{\max})}[i]$.
The final transmitted signal is $x_{k,\ell}^{(\mathrm{ICF})}[i]$, which satisfies the peak-power constraint with a certain threshold for out-of-band distortion. The flow diagram of the whole procedure can be found in Fig. 2.
IV Simulations
IV-A Simulation Settings
[…]
IV-B Simulation Results and Discussions
All curves in this section are smoothed for readability.
Peak-to-average power ratio is shown in Fig. 3. Without clipping, the PAPR is high (about $35\text{\,}\mathrm{dB}$), while after clipping and filtering, the PAPR remains close to $15\text{\,}\mathrm{dB}$. Note that the results without clipping and before clipping are different since PAPR depends on the actual values to be transmitted and clipping and filtering procedure changes the actual values to approximate the average of gradient (5). We also observe that single-carrier transmission can exhibit higher PAPR than multi-carrier OFDM transmission, which is unusual for digital modulation schemes. This stems from the nature of amplitude modulation (for the single-carrier case), where the power is directly related to the actual values to be transmitted, which can actually go up to “infinity” if the gradient values can take arbitrarily large values. The decrease in PAPR in the single-carrier case w.r.t. communication rounds may be due to that the model has been well trained, so the gradient values divergence is smaller.
The approximation error for (5), i.e. the MSE in (8) and (16), is shown in Fig. 4. The true squared error (TSE) is defined as $\frac{1}{N}\sum_{n=1}^{N}\|\tilde{g}[n]-\overline{g}[n]\|^{2}$. Clipping increases the TSE relative to the non-clipping baseline, and the analytical MSE underestimates the TSE. Interestingly, for multi-carrier OFDM, the post-clipping TSE is larger when the noise power is lower. The reason is that, at low noise, the scaling factors $\alpha_{n}$ take larger values, therefore amplifying the in-band distortion caused by iterative clipping-and-filtering, resulting in higher error.
[…]
(a) Single-carrier.
(b) Multi-carrier.
Figure 3: Peak-to-average power ratio (PAPR) w.r.t. communication rounds.
(a) Single-carrier
(b) Multi-carrier.
Figure 4: Error of the approximated average gradient w.r.t. the actual average gradient. TSE denotes the True Square Error; Estimated MSE is given by (8) and (16).
(a) Single-carrier
(b) Multi-carrier.
Figure 5: Test Accuracy w.r.t. communication rounds.
(a) Single-carrier
(b) Multi-carrier.
Figure 6: Loss of test accuracy by imposing the peak power constraint (after clipping and filtering).
V Conclusion
This work highlights an overlooked practical issue in over-the-air computation-federated learning: the instantaneous peak power constraint imposed by the nonlinearity in power amplifiers. This is an issue in over-the-air computation-federated learning since it uses analog transmission and some works consider multi-carrier OFDM transmissions. We applied the classic PAPR reduction methods of iterative amplitude clipping and frequency-domain filtering. Enforcing peak power constraint can indeed degrade aggregation and final accuracy, in some cases, the effects are marginal, and in some other cases (multi-carrier with low noise level), cause divergence. Counterintuitively, the impact is stronger at lower noise for multi-carrier OFDM because the scaling factor from optimization is higher and therefore the in-band distortion is too much amplified. This work should motivate further research in peak power constraint-aware power control and better PAPR reduction techniques for single-carrier and multi-carrier over-the-air computation-federated learning systems.

### v2

@@
 Federated learning has been considered a promising privacy preserving distributed edge learning framework.
-Over-the-air computation technique leveraging analog transmission enables the aggregation of local updates directly over-the-air by exploiting the superposition properties of wireless multiple-access channel, thereby drastically reducing the communication bottleneck issues of federated learning compared with digital transmission schemes.
-This work points out that existing over-the-air computation-federated learning overlooks a key practical constraint, the instantaneous peak-power constraints imposed by the non-linearity of radio-frequency power amplifiers.
-We present and analyze the effect of the classic method to deal with this issue, amplitude clipping combined with filtering.
-We investigate the effect of instantaneous peak-power constraints in over-the-air computation-federated learning for both single-carrier and multi-carrier orthogonal frequency-division multiplexing systems.
-We highlight the specificity of over-the-air computation-federated learning: the samples depend on the gradient value distribution, leading to a higher peak-to-average power ratio than that observed for uniformly distributed signals.
-Simulation results demonstrate that, in practical settings, the instantaneous transmit power regularly exceeds the power-amplifier limit; however, by applying clipping and filtering, the FL performance can be degraded.
+Over-the-air computation leveraging analog transmission enables the aggregation of local updates directly over-the-air by exploiting the superposition properties of wireless multiple-access channels, thereby alleviating the communication bottleneck issues of federated learning compared with digital transmission schemes.
+This work points out that existing over-the-air computation-federated learning overlooks a key practical constraint, the instantaneous peak-power constraints due to the non-linearity of radio-frequency power amplifiers.
+Operating directly in non-linear region causes in-band and out-of-band distortions.
+We present and analyze the effect of the default method that limits the signal’s peak power and out-of-band distortions, iterative amplitude clipping combined with filtering.
+We investigate the effect of imposing instantaneous peak-power constraints in over-the-air computation-federated learning for both single-carrier and multi-carrier orthogonal frequency-division multiplexing systems.
+Simulation results demonstrate that, in practical settings, the instantaneous transmit power in AirComp-FL regularly exceeds the power-amplifier linearity limit.
+As the first work of this line of research, it is essential to evaluate if this is an actual problem that has an impact on FL performance.
+We therefore apply the classic method of iterative clipping and filtering, and show that the FL performance degrades more or less depending on the scenarios.
 The degradation becomes pronounced especially in multi-carrier OFDM systems due to the in-band distortions caused by clipping and filtering.
 Index Terms:
-federated learning, over-the-air computation, peak-to-average, peak power constraint.
+federated learning, over-the-air computation (AirComp), peak-to-average power ratio (PAPR), peak power constraint.
+publicationid: pubid: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
 I Introduction
-An unprecedented volume of data is now generated at the wireless edge.
-Leveraging these data to train machine learning models at a centralized server is challenged by privacy concerns and limited uplink bandwidth requirements.
-Federated learning addresses both challenges by keeping raw data on devices and only exchanging model updates [1].
-Federated learning consists of a collaborative learning framework, in which repeatedly, each edge device trains on its own data for certain iterations and then uploads the model or the gradients to a central server for aggregation.
-However, the digital transmission of large-scale model updates still creates a communication bottleneck, especially as the number of participants scales [2].
-As more and more end devices are wirelessly connected, considering federated learning within wireless networks is essential [3].
-Over-the-air computation technique proposed in [4, 5] is investigated as a means to counter the communication bottleneck issue of federated learning [6, 7, 8].
-By exploiting the waveform superposition of the wireless multiple-access channel, a server can directly aggregate local updates “over the air,” enabling scalability with the number of devices to participate in each communication round.
-One of the central design challenges in over-the-air computation-federated learning is power control for signal power or amplitude alignment.
-The ideal aim is to obtain equally average received power from user equipments at the base station so that a summation operation can be performed.
-Under average power constraints, channel inversion and threshold-based policies have been shown to minimize aggregation mean-squared error: user equipments with weak channels transmit at their maximum power budget, while user equipments above a threshold invert their channels [9, 10, 11].
-However, the above designs ignore the instantaneous peak-power limitations due to the non-linearity of power amplifiers from a certain power level.
-Operating in the nonlinear region causes signal distortion and receiver misalignment.
-This is a common issue for orthogonal frequency-division multiplexing, whose peak-to-average power ratio is known to be high.
-While recent works have advanced over-the-air computation in orthogonal frequency-division multiplexing systems [12, 13, 14], they typically optimize mean-squared error under average-power budgets and imperfect channel state information but do not quantify or enforce instantaneous peak-power constraints.
-This paper studies over-the-air computation-federated learning under both average and instantaneous peak-power constraints for single-carrier transmission and multi-carrier orthogonal frequency-division multiplexing systems.
-We characterize the induced aggregation error when peaks are constrained, and evaluate the effect of practical mitigation mechanisms.
-The main contributions are summarized as follows:
-  • We highlight the practical limitations of peak power constraint in a practical system due to the non-linear region of power amplifiers.
-  • We present and evaluate the most common method to handle the peak power constraints: amplitude clipping that repeatedly applies clipping, i.e., clip the time-domain peak power value that is set by the nonlinearity of the power amplifier, and filtering the out-of-band distortion.
-Implementation details for both multiple-access schemes are provided.
-  • Empirical evaluation and design insights.
-Evaluating under LeNet trained on CIFAR-10, we quantify peak-to-average power ratio, aggregation mean-squared error, and learning accuracy.
-We observe substantially high peak-to-average power ratio of a over-the-air computation-federated learning signal, either in single-carrier or multi-carrier OFDM.
-Enforcing peak constraints via clipping yields accuracy degradation and may cause model divergence.
-The remaining sections are organized as follows: Section II introduces the FL-OTA system models for single-carrier and multi-carrier OFDM and the associated average power allocation based on MSE minimization.
-Section III formalizes the instantaneous peak-power constraint and presents the clipping mechanisms.
-Section IV presents the simulation settings and results.
+Federated learning, as a distributed edge learning paradigm, addresses privacy and security concerns of deep learning by keeping raw data on devices and only transmitting model updates for aggregation [1].
+However, digital transmission of large-scale model updates of massive edge devices creates a communication bottleneck that does not scale well with the number of edge devices [2, 3].
+Over-the-air computation [4, 5] counters this bottleneck in federated learning [6, 7, 8] by exploiting waveform superposition of the wireless multiple-access channel to aggregate local updates directly over the air.
+A central design challenge is power control for amplitude alignment: under average power constraints, threshold-based policies minimizing aggregation mean-squared error have been developed, where devices with weak channels transmit at maximum power and stronger-channel devices invert their channels [9, 10, 11].
+However, the above designs ignore the instantaneous peak-power limitations due to the non-linearity of power amplifiers as illustrated in Fig.
+1.
+Given a certain average input power $P_{\mathrm{avg}}^{(in)}$, the actual operating region spans up to the maximum peak instantaneous power obtained by multiplying $P_{\mathrm{avg}}^{(in)}$ with the signal peak-to-average power ratio.
+Operating in the nonlinear region causes in-band distortions that corrupt over-the-air computation aggregation and out-of-band emissions that must be strictly limited, since they may interfere with adjacent frequency bands.
+In over-the-air computation, signal amplitude depends directly on gradient value distributions and may thus encounter high peak-to-average power ratio; multi-carrier orthogonal frequency-division multiplexing further inherits its well-known high-peak-to-average power ratio problem.
+While recent works advance over-the-air computation-orthogonal frequency-division multiplexing [12, 13, 14], they optimize mean-squared error under average-power budgets but do not quantify or enforce instantaneous peak-power constraints.
+This paper studies over-the-air computation-federated learning under peak-power constraints for single-carrier and multi-carrier orthogonal frequency-division multiplexing systems.
+The main contributions are:
+  • We highlight the instantaneous peak-power constraint imposed by power amplifier non-linearity in over-the-air computation-federated learning for both single-carrier and multi-carrier orthogonal frequency-division multiplexing.
+  • As the first work to highlight this practical issue of high peak-to-average power ratio in the context of AirComp-FL, it is essential to confirm that this is an actual issue that can have an impact on FL performance.
+We therefore propose a detailed characterization of the problem and evaluate the effect of the most classic solution for addressing this issue, iterative clipping-and-filtering, with implementation details for both transmission schemes.
+  • Empirical evaluation under LeNet/CIFAR-10: both single-carrier and multi-carrier over-the-air computation-federated learning exhibit high peak-to-average power ratio; clipping degrades accuracy and can cause model divergence.
+Figure 1: Typical input and output power characteristics curve for a power amplifier.
 II System Model
 II-A Federated Learning
-Consider a single-antenna base station serving a set $\mathcal{K}\triangleq\{1,\ldots,K\}$ of $K$ user equipments, where federated learning is performed.
-A supervised learning task is considered for FL.
-User equipment $k$ holds a local dataset $\mathcal{D}_{k}$ of size $D_{k}$.
-The local empirical loss function can be written as
-$F_{k}(\bm{{w}})=\frac{1}{D_{k}}\sum_{x\in\mathcal{D}_{k}}\ell(\bm{{w}};x),$ (1)
-where $\ell(\bm{{w}};x)$ denotes the loss of the prediction of the sample $x\in\mathcal{D}_{k}$ evaluated with model weights $\bm{{w}}\in\mathbb{R}^{N}$.
-Federated learning has the objective of minimizing the empirical global loss function:
-$\min_{\bm{{w}}\in\mathbb{R}^{N}}\sum_{k\in\mathcal{K}}D_{k}F_{k}(\bm{{w}}).$ (2)
-Equal size of local dataset is assumed, i.e. $D_{k}=D$ for any $k\in\mathcal{K}$. Federated learning consists of multiple communication rounds where $t$-th communication round is described as follows:
+Consider a single-antenna base station serving $K$ user equipments, where federated learning is performed.
+User equipment $k$ holds local dataset $\mathcal{D}_{k}$ with local empirical loss $F_{k}(\bm{{w}})=\frac{1}{D_{k}}\sum_{x\in\mathcal{D}_{k}}\ell(\bm{{w}};x)$, where $\ell(\bm{{w}};x)$ is the per-sample loss on $x$ with model weights $\bm{{w}}\in\mathbb{R}^{N}$.
+Federated learning minimizes $\min_{\bm{{w}}}\sum_{k\in\mathcal{K}}D_{k}F_{k}(\bm{{w}})$, assuming here equal dataset sizes $D_{k}=D$ [6].
+Each communication round $t$ proceeds as follows:
   1. 1.
-    Base station broadcasts the global model from previous communication round $\bm{{w}}^{(t-1)}$ to all user equipments.
-This step is assumed without issue since base station has sufficient transmit power.
+    Base station broadcasts global model $\bm{{w}}^{(t-1)}$ to all user equipments.
   2. 2.
-    Each user equipment $k\in\mathcal{K}$ performs a mini-batch stochastic gradient descent step based on $w^{(t-1)}$ to minimize the local empirical loss $F_{k}$ over its local dataset $\mathcal{D}_{k}$.
-The resulting gradient vector is denoted as $\overline{\bm{{g}}}_{k}^{(t)}$.
+    Each user equipment $k$ computes local gradient $\overline{\bm{{g}}}_{k}^{(t)}$ via mini-batch SGD on $\mathcal{D}_{k}$.
   3. 3.
-    When all user equipments are ready for uplink transmission, a pilot signal is sent orthogonally from each user equipment to base station (assuming sufficient resources for short pilot signals).
-We assume perfect channel state information at the base station as commonly assumed in the literature [8, 6].
-With the channel state of each user equipment, base station determines the transmit power that each user equipment should use for over-the-air computation, and sends the power information to each user equipment.
+    User equipments send pilots; base station estimates channels and feeds back optimal transmit powers (perfect channel state information assumed [8, 6]).
   4. 4.
-    All user equipments apply the same normalization that transforms $\overline{\bm{{g}}}_{k}^{(t)}$ to $\bm{{g}}_{k}^{(t)}$ so that $\operatorname{\mathbb{E}}[\bm{{g}}_{k}^{(t)}]=0$ and $\operatorname{\mathbb{E}}[\|\bm{{g}}_{k}^{(t)}\|^{2}_{2}]=N$ [8, 15]:
+    All user equipments apply a shared normalization [8, 15] so that $\operatorname{\mathbb{E}}[\bm{{g}}_{k}^{(t)}]=0$, $\operatorname{\mathbb{E}}[\|\bm{{g}}_{k}^{(t)}\|^{2}_{2}]=N$:
     $\bm{{g}}_{k}^{(t)}=\frac{\overline{\bm{{g}}}_{k}^{(t)}-\mu}{\Gamma},\vskip-2.84544pt$ (1)
-    where the normalizing mean $\mu^{(t)}=\frac{1}{|\mathcal{S}_{c}|}\sum_{k\in\mathcal{S}_{c}}\mu_{k}^{(t)}$ and standard deviation $\Gamma^{(t)}=\frac{1}{|\mathcal{S}_{c}|}\sum_{k\in\mathcal{S}_{c}}\Gamma_{k}^{(t)}$ are obtained by averaging the statistics of each participating clients:
-    $\mu_{k}^{(t)}=\frac{1}{m}\sum_{i=1}^{d}\overline{g}_{k}^{(t)}[i],\quad\Gamma_{k}^{(t)}=\sqrt{\frac{1}{m}\sum_{i=1}^{d}\overline{g}_{k}^{(t)2}[i]-\mu_{k}^{(t)2}},$ (4)
-    where $\overline{\bm{{g}}}_{k}^{(t)}=(\overline{g}_{k}^{(t)}[i])_{i=1,\ldots,N}$ the gradient vector for end-device $k$.
-Note that $\mu^{(t)}$ and $\Gamma^{(t)}$ have to be uniform across all end-devices due to the nature of AirComp and must be estimated before the data communication.
+    with $\mu^{(t)}=\frac{1}{|\mathcal{S}_{c}|}\!\sum_{k\in\mathcal{S}_{c}}\!\mu_{k}^{(t)}$, $\Gamma^{(t)}=\frac{1}{|\mathcal{S}_{c}|}\!\sum_{k\in\mathcal{S}_{c}}\!\Gamma_{k}^{(t)}$, where $\mu_{k}^{(t)}\!=\!\frac{1}{N}\!\sum_{i}\overline{g}_{k}^{(t)}[i]$ and $\Gamma_{k}^{(t)}\!=\!\sqrt{\frac{1}{N}\!\sum_{i}\overline{g}_{k}^{(t)2}[i]-\mu_{k}^{(t)2}}$.
   5. 5.
-    Each then sends the normalized gradient $\bm{{g}}_{k}^{(t)}$ via analog signal with the power control information given by base station.
-The results via over-the-air computation at the base station gives the approximation of the gradient average $\tilde{\bm{{g}}}^{(t)}$:
+    User equipments transmit $\bm{{g}}_{k}^{(t)}$ via analog signal; base station recovers the gradient average via AirComp
     $\tilde{\bm{{g}}}^{(t)}\approx\overline{\bm{{g}}}^{(t)}\triangleq\frac{1}{K}\sum_{k\in\mathcal{K}}\overline{\bm{{g}}}_{k}^{(t)},\vskip-2.84544pt$ (2)
-    (details will be shown in later section), and update the new global model $\bm{{w}}^{(t)}=\bm{{w}}^{(t-1)}-\eta\overline{\bm{{g}}}^{(t)}$ with $\eta>0$ the learning rate and $\overline{\bm{{g}}}^{(t)}$ the de-normalized gradients.
-In the following sections, we focus on step 5 that consists of the gradient transmission via over-the-air computation techniques in single-carrier and multi-carrier OFDM-based systems.
-For ease of notation, the index of communication round $t$ is removed from all notations, since the principle is identical given any channel states and gradient vectors.
+    and updates $\bm{{w}}^{(t)}=\bm{{w}}^{(t-1)}-\eta\overline{\bm{{g}}}^{(t)}$.
+For ease of notation, the round index $t$ is dropped below.
 II-B Single-Carrier over-the-air computation
@@
 $=\frac{\Gamma^{2}N}{K^{2}}\Big(\sum_{k\in\mathcal{K}}(h_{k}\sqrt{\alpha p_{k}}-1)^{2}+\alpha\sigma^{2}\Big).$
-The equation can be derived by independence of zero-mean Gaussian noise with $\bm{{g}}_{k}$ and that $\operatorname{\mathbb{E}}[\|\bm{{g}}_{k}\|_{2}^{2}]=N$ and $\operatorname{\mathbb{E}}[\bm{{g}}_{k}]=0$.
-A similar expression with a factor $N+1$ can also be obtained without such assumptions on the gradient vector, which can be obtained via the Cauchy-Schwarz inequality as shown in [16].
+The derivation uses independence of zero-mean Gaussian noise with $\bm{{g}}_{k}$, $\operatorname{\mathbb{E}}[\|\bm{{g}}_{k}\|_{2}^{2}]=N$, and $\operatorname{\mathbb{E}}[\bm{{g}}_{k}]=0$.
 To optimize the MSE, we optimize the term in the outer bracket subject to the average power constraint $P_{\mathrm{avg},\max}$. The resulting optimization problem is given by
@@
 II-C1 Transmission of gradient
-In multi-carrier OFDM systems, individual values of $\bm{{g}}_{k}$ are modulated on subchannels in the frequency domain by assigning one value of $\bm{{g}}_{k}$ to one subchannel.
-Since the number of subchannels $M$ within the system bandwidth $B$ is typically less than the size of model parameter $N$, transmission spans multiple OFDM symbols.
+In multi-carrier OFDM systems, individual values of $\bm{{g}}_{k}$ are modulated on subcarriers in the frequency domain by assigning one value of $\bm{{g}}_{k}$ to one subcarrier.
+Since the number of subcarriers $M$ within the system bandwidth $B$ is typically less than the size of model parameter $N$, transmission spans multiple OFDM symbols.
 Let the OFDM symbol index be $\ell\in\mathcal{S}\triangleq\{0,\ldots,L-1\}$ with $L=\lceil\frac{N}{M}\rceil$.
@@
 Zero values are transmitted for $n\geq N$.
-The $i$-th transmitted OFDM-symbol discrete sample in the time domain of the $k$-th user equipment and $\ell$-th OFDM-symbol is given by its IDFT:
+The $i$-th time-domain sample of the $k$-th user equipment’s $\ell$-th OFDM symbol is:
 $(\forall i\in\mathcal{M})\quad s_{k,\ell}[i]=\sum_{m=0}^{M-1}g_{k}[\ell M+m]b_{k,\ell M+m}e^{j\frac{2\pi mi}{M}},\vskip-2.84544pt$ (9)
-where $b_{k,\ell M+m}\in\mathbb{C}$ scales the transmitted power to mitigate the effect of the channel.
-The received frequency domain signal is given by the transmitted signal $s_{k,\ell}[i]$ convolved with the channel impulse response $h_{k,\ell}[i]$ with the length $C+1$.
-We suppose that each device transmits its gradient perfectly synchronized, so that the received signal at the sample $i$, such that $C\leq i\leq M-1$ becomes
-$y[i]$ $=\sum_{k\in\mathcal{K}}s_{k,\ell}[i]\circledast h_{k,\ell}[i]+n[i],$ (13)
-$\hskip-22.76228pt=\sum_{k\in\mathcal{K}}\!\Big(\!\sum_{m=0}^{M-1}\!g_{k}[\ell M\!+\!m]b_{k,\ell M+m}e^{j\frac{2\pi mi}{M}}\!\Big)\!\circledast\!h_{k,\ell}[i]\!+\!n[i],$
-[…]
-$Y_{\ell M+m}$ $=\mathrm{DFT}\{\{y[i]\}_{i\in\mathcal{M}}\}[m]$ (14)
-$\hskip-25.6073pt=\frac{1}{M}\sum_{k\in\mathcal{K}}\!\sum_{i=0}^{M-1}\!\Big(\!\sum_{m^{\prime}=0}^{M-1}\!g_{k}[\ell M\!+\!m^{\prime}]b_{k,\ell M+m^{\prime}}e^{j\frac{2\pi m^{\prime}i}{M}}\!\circledast\!h_{k,\ell}[i]$
-$\quad\quad+n[i]\Big)e^{-j\frac{2\pi mi}{M}}$
-$\hskip-25.6073pt=\sum_{k=1}^{K}g_{k}[\ell M+m]b_{k,\ell M+m}H_{k,\ell M+m}+N[m],$
-where $H_{k,\ell M+m}=|H_{k,\ell M+m}|e^{j\phi_{k,\ell M+m}}$ is the frequency response of $h_{k,\ell}[i]$ and $N[m]\sim\mathcal{C}\mathcal{N}(0,\sigma^{2})$ is the additive white Gaussian noise for each sub channel since the Fourier transform of a Gaussian distribution is also a Gaussian distribution with the same statistical properties.
-Due to the fact that each sub channel is associated with a different channel $H_{k,\ell M+m}$ in OFDM, the $n$-th gradient value can therefore be recovered by:
-$\tilde{g}[n]$ $=\frac{\Gamma\sqrt{\alpha_{n}}Y_{n}}{K}+\mu,$ (15)
-$\hskip-14.22636pt=\frac{\Gamma\sqrt{\alpha_{n}}}{K}\sum_{k\in\mathcal{K}}|H_{k,n}|\sqrt{p_{k,n}}g_{k}[n]+\frac{\Gamma\sqrt{\alpha_{n}}}{K}N[n]+\mu,$
-where $b_{k,\ell M+m}$ in (14) is substituted as $b_{k,\ell M+m}=\sqrt{p_{k,\ell M+m}}e^{-j\phi_{k,\ell M+m}}$, with $p_{k,\ell M+m}$ the transmit power of UE $k$ at the subcarrier $m$ when transmitting the $\ell$-th symbol, and hence already includes the phase offset caused by the channel and $\alpha_{n}>0$ is the denoising factor for the corresponding subcarrier.
+where $b_{k,\ell M+m}\in\mathbb{C}$ performs channel pre-equalization.
+Assuming synchronized transmission and cyclic-prefix length exceeding the channel delay spread, applying DFT at the base station yields for $m\in\mathcal{M}$:
+$Y_{\ell M+m}=\sum_{k=1}^{K}g_{k}[\ell M+m]\,b_{k,\ell M+m}\,H_{k,\ell M+m}+N[m],\vskip-2.84544pt$ (10)
+where $H_{k,n}=|H_{k,n}|e^{j\phi_{k,n}}$ is the sub-channel response and $N[m]\sim\mathcal{C}\mathcal{N}(0,\sigma^{2})$.
+Setting $b_{k,n}=\sqrt{p_{k,n}}e^{-j\phi_{k,n}}$, the $n$-th gradient value is recovered as:
+$\tilde{g}[n]=\frac{\Gamma\sqrt{\alpha_{n}}}{K}\sum_{k\in\mathcal{K}}|H_{k,n}|\sqrt{p_{k,n}}\,g_{k}[n]+\frac{\Gamma\sqrt{\alpha_{n}}}{K}N[n]+\mu,\vskip-5.69046pt$ (11)
+with $p_{k,n}$ the transmit power of UE $k$ at subcarrier $n$ and $\alpha_{n}>0$ the denoising factor.
 II-C2 Power Allocation
@@
 $\mathrm{MSE}_{n}$ $=\operatorname{\mathbb{E}}[\|\tilde{g}[n]-\overline{g}[n]\|^{2}]$ (12)
-$=\frac{\Gamma^{2}}{K^{2}}\Big(\sum_{k\in\mathcal{K}}(|H_{k,n}|\sqrt{\alpha_{n}p_{k,n}}-1)^{2}+\alpha_{n}\sigma^{2}\Big),$
+$=\frac{\Gamma^{2}}{K^{2}}\Big(\sum_{k\in\mathcal{K}}(|H_{k,n}|\sqrt{\alpha_{n}p_{k,n}}-1)^{2}+\alpha_{n}\sigma^{2}\Big).$
 For any OFDM-symbol $\ell$, the optimal power allocation for gradient values of indices $n=\ell M+m$ with $m=0,\ldots,M-1$ can be given as:
@@
 This problem has been analyzed in [13] that requires a subgradient solver and partly in [14] that considers an equal $\alpha$ for all subcarriers.
-Due to the short time frame that the power needs to be allocated, only low-complexity solution is practically feasible.
-Since transmit power at the UE side is typically low, we pick up the classical empirical observations that the power allocation for uplink multi-carrier OFDM has a marginal effect on the performance [17, 18].
-Assuming equal power allocated to each subcarrier for each UE, i.e., replacing the constraint (17b) by
-$(\forall k\in\mathcal{K})\ (\forall m\in\mathcal{M})\ p_{k,\ell M+m}\leq\frac{P_{\mathrm{avg},\max}}{M}.\vskip-4.26773pt$ (18)
-The problem can therefore be decoupled for each subcarrier $m$, and the optimal solution for each subcarrier $m$ given a symbol $\ell$, i.e., given $(H_{k,\ell M+m})_{k\in\mathcal{K}}$, $P_{\mathrm{avg},\max}/M$ as the maximum average transmit power, and $\sigma^{2}$ as the noise power level, can be calculated again with Proposition 1.
+Due to the short allocation time and the marginal effect of per-subcarrier power control in uplink OFDM [17, 18], we assume equal max power per subcarrier, replacing (13b) by $p_{k,\ell M+m}\leq P_{\mathrm{avg},\max}/M$.
+The problem then decouples per subcarrier $m$, each solved via Proposition 1 with budget $P_{\mathrm{avg},\max}/M$.
 III Instantaneous peak power constraint
-Figure 1: Typical input and output power characteristics curve for a power amplifier.
-Practical radio-frequency power amplifiers operate linearly only up to certain input power [19, 20].
-Given a certain average power as input $P_{\mathrm{avg}}^{(in)}$, the actual operating region would be between zero and the maximum peak instantaneous input power which can be obtained by multiplying $P_{\mathrm{avg}}^{(in)}$ with peak-to-average power ratio (addition in dB domain), as illustrated in Fig.
-1.
-Operating in nonlinear region leads to signal distortions, resulting in strong in-band and out-of-band distortions (details see later), which is to be avoided.
-Note that all previously determined powers are actually output powers, but can be easily converted to the input power by a one-to-one relationship.
+As presented in the introduction, practical radio-frequency power amplifiers operate linearly only up to certain input power [19, 20].
+Operating in linear region would result in in-band and out-of-band distortions.
+Note that all previously determined powers are actually output powers from the power amplifier, but can be easily converted to the input power by a one-to-one relationship.
 We denote $p_{k}^{(in)}$ and $b_{k\ell M+m}^{(in)}$ respectively the corresponding input power of $p_{k}$ and $b_{k\ell M+m}$.
@@
 $\mathrm{PAPR}_{k,\ell}^{\mathrm{ideal}}=\frac{\max_{t}|s_{k,\ell}(t)|^{2}}{\frac{1}{T}\int_{0}^{T}|s_{k,\ell}(t)|^{2}dt}.$ (17)
-In practice, only digital discrete samples can be employed to estimate PAPR. It has been shown that samples sampled at Nyquist rate inaccurately captures PAPR.
-Oversampling: To accurately capture PAPR in the time-domain OFDM signal, oversampling is essential.
-An oversampling factor $L_{\mathrm{os}}$ (typically $L_{\mathrm{os}}=4$ is sufficient [21]) is applied by zero-padding the frequency-domain signal before the IDFT.
-The oversampled time-domain signal is generated as:
-$s_{k,\ell}^{(\mathrm{os})}[i]=\!\!\!\sum_{m=0}^{L_{\mathrm{os}}M-1}\!\!G_{k,\ell}[m]e^{j\frac{2\pi mi}{L_{\mathrm{os}}M}},\quad i\!=\!0,\ldots,L_{\mathrm{os}}M\!-\!1,\vskip-2.84544pt$ (23)
-where the zero-padded frequency-domain vector is defined as:
-$\hskip-7.11317ptG_{k,\ell}[m]\!=\!\begin{cases}g_{k}[\ell M\!+\!m]b_{k,\ell M+m}^{(in)},&0\!\leq\!m\!<\!M,\\ 0,&M\!\leq\!m\!<L_{\mathrm{os}}M.\end{cases}\vskip-2.84544pt$ (24)
-The per-symbol PAPR can therefore be estimated as
+In practice, Nyquist-rate samples underestimate the true PAPR.
+An oversampling factor $L_{\mathrm{os}}=4$ [21] is applied by zero-padding the frequency-domain signal ($G_{k,\ell}[m]=g_{k}[\ell M+m]b_{k,\ell M+m}^{(in)}$ for $0\leq m<M$, zero otherwise) before the IDFT:
+$s_{k,\ell}^{(\mathrm{os})}[i]=\!\!\!\sum_{m=0}^{L_{\mathrm{os}}M-1}\!\!G_{k,\ell}[m]e^{j\frac{2\pi mi}{L_{\mathrm{os}}M}},\quad i\!=\!0,\ldots,L_{\mathrm{os}}M\!-\!1.\vskip-2.84544pt$ (18)
+The per-symbol PAPR is then estimated as
 $\mathrm{PAPR}_{k,\ell}=\frac{\max_{i}|s_{k,\ell}^{(\mathrm{os})}[i]|^{2}}{\frac{1}{L_{\mathrm{os}}M}\sum_{i=0}^{L_{\mathrm{os}}M-1}|s_{k,\ell}^{(\mathrm{os})}[i]|^{2}}.\vskip-4.26773pt$ (19)
 III-B Clipping with Oversampling
-The commonly used technique to deal with the high peak-to-average power ratio issue is applying the procedure of iterative clipping-and-filtering.
-Since clipping results in in-band and out-of-band distortions in multi-carrier OFDM [19, 20].
-A low-pass filter is applied afterwards to limit out-of-band distortion, which may result in the rise of PAPR, necessitating repeated such procedures.
+The most classic technique without changing the modulation schemes (that are necessary for over-the-air computation) to deal with the high peak-to-average power ratio issue is applying the procedure of iterative clipping-and-filtering [19].
+Since clipping results in in-band and out-of-band distortions in multi-carrier OFDM [20], a low-pass filter is applied afterwards to limit out-of-band distortion, which may result in the rise of PAPR, necessitating repeating such procedures.
 Amplitude Clipping
@@
   • In-band distortion: distortion on the considered subcarriers $m\in\mathcal{M}$, which degrades the aggregated gradients.
-  • Out-of-band emissions: spectral regrowth on the zero-padded subcarriers $M\leq m<L_{\mathrm{os}}M$ due to clipping, which causes interference to adjacent channels.
+  • Out-of-band emissions: spectral regrowth on the zero-padded subcarriers $M\leq m<L_{\mathrm{os}}M$ due to clipping, which is to be strictly limited since it may cause interference to adjacent frequency bands.
 Filtering
-To suppress out-of-band emissions, a low-pass filteris applied. Define a rectangular filter:
-$H_{\mathrm{filter}}[m]=\begin{cases}1,&0\leq m<M,\\ 0,&M\leq m<L_{\mathrm{os}}M.\end{cases}\vskip-2.84544pt$ (29)
-The filtered frequency-domain signal is:
-$X_{k,\ell}^{(\mathrm{filt})}[m]=X_{k,\ell}[m]\cdot H_{\mathrm{filter}}[m].\vskip-2.84544pt$ (30)
-Converting back to the time domain yields the filtered signal:
-$x_{k,\ell}^{(\mathrm{filt})}[i]=\mathrm{IDFT}\bigl\{X_{k,\ell}^{(\mathrm{filt})}[m]\bigr\}[i].\vskip-4.26773pt$ (31)
-However, filtering may reintroduce peak regrowth, as it restores some of the constructive interference among subcarriers that clipping had disrupted [19].
+To suppress out-of-band emissions, a rectangular low-pass filter ($H_{\mathrm{filter}}[m]=1$ for $0\leq m<M$, else 0) is applied in the frequency domain and converted back via IDFT:
+$x_{k,\ell}^{(\mathrm{filt})}[i]=\mathrm{IDFT}\bigl\{X_{k,\ell}[m]\cdot H_{\mathrm{filter}}[m]\bigr\}[i].\vskip-2.84544pt$ (23)
+However, filtering may reintroduce peak regrowth [19].
 Iterative Clipping and Filtering
-To jointly control PAPR and out-of-band emissions, an iterative clipping-and-filtering procedure is commonly employed [22]. The algorithm proceeds as follows for each symbol $\ell$:
+To jointly control PAPR and out-of-band emissions, an iterative clipping-and-filtering procedure is commonly employed [22]:
   1. 1.
-    Initialize: $x_{k,\ell}^{(0)}[i]=s_{k,\ell}^{(\mathrm{os})}[i]$; $j=1$.
+    Set $x_{k,\ell}^{(0)}[i]=s_{k,\ell}^{(\mathrm{os})}[i]$.
   2. 2.
-    While $x_{k,\ell}^{(j,\mathrm{clip})}[i]$ has out-of-band distortion higher than a threshold:
-    1. (a)
-      Clip: $x_{k,\ell}^{(j,\mathrm{clip})}[i]=\mathcal{C}_{A_{\max}}\bigl(x_{k,\ell}^{(j-1)}[i]\bigr)$.
-    2. (b)
-      DFT: $X_{k,\ell}^{(j)}[m]=\mathrm{DFT}\bigl\{x_{k,\ell}^{(j,\mathrm{clip})}[i]\bigr\}[m]$.
-    3. (c)
-      Filter: $X_{k,\ell}^{(j,\mathrm{filt})}[m]=X_{k,\ell}^{(j)}[m]\cdot H_{\mathrm{filter}}[m]$.
-    4. (d)
-      IDFT: $x_{k,\ell}^{(j)}[i]=\mathrm{IDFT}\bigl\{X_{k,\ell}^{(j,\mathrm{filt})}[m]\bigr\}[i]$.
+    Repeat until out-of-band distortion is below threshold: clip $x^{(j-1)}$ to $A_{\max}$; apply DFT; apply $H_{\mathrm{filter}}$; apply IDFT to obtain $x^{(j)}$.
   3. 3.
-    Output: $x_{k,\ell}^{(\mathrm{ICF})}[i]=x_{k,\ell}^{(J_{\max})}[i]$.
-The final transmitted signal is $x_{k,\ell}^{(\mathrm{ICF})}[i]$, which satisfies the peak-power constraint with a certain threshold for out-of-band distortion.
-The flow diagram of the whole procedure can be found in Fig.
-2.
+    Output $x_{k,\ell}^{(\mathrm{ICF})}[i]=x_{k,\ell}^{(J_{\max})}[i]$.
+The flow diagram is shown in Fig. 2.
 IV Simulations
 IV-A Simulation Settings
-[…]
+In the simulations the LeNet convolutional neural network with $62006$ trainable parameters is trained on the Cifar-10 dataset for $500$ communication rounds, where the data is independent and identically distributed among $K=40$ user equipments.
+In each communication round all user equipments participate and are uniformly distributed on a $100$ meter radius disk around the base station.
+As a large-scale fading model, the free-space-path loss is used, which is combined with Rayleigh fading as a small-scale fading model.
+The constraint of the maximum average power $P_{\mathrm{avg},\max}$ of user equipments is chosen to be $23\text{\,}\mathrm{dBm}$, while the constraint of the the maximum instantaneous power $P_{\mathrm{inst},\max}$ is chosen to be $26\text{\,}\mathrm{dBm}$.
+In multi-carrier OFDM transmission, $M=32$ subcarriers are used in order to reduce OFDM complexity, where each subcarrier transmits with a bandwidth of $60\text{\,}\mathrm{kHz}$.
+Clipping and filtering iterate until out-of-band emissions are below $10\text{\,}\mathrm{dBm}$.
+The learning rate is set at 1 and the local batch size at 256.
 IV-B Simulation Results and Discussions
@@
 The decrease in PAPR in the single-carrier case w.r.t.
-communication rounds may be due to that the model has been well trained, so the gradient values divergence is smaller.
+communication rounds may be due to that the model has been well trained, so the gradient value variance is smaller.
 The approximation error for (2), i.e.
@@
 The reason is that, at low noise, the scaling factors $\alpha_{n}$ take larger values, therefore amplifying the in-band distortion caused by iterative clipping-and-filtering, resulting in higher error.
-[…]
+The results on FL test accuracy are shown in Fig.
+5.
+For single-carrier (Fig.
+5a), we observe consistent but small gaps between clipped and unclipped cases, with higher noise yielding lower accuracy overall.
+A noticeable accuracy drop appears at around 50 communication rounds for the clipped case.
+For multi-carrier OFDM (Fig.
+5b), as predicted by the TSE, the clipped case performs worse at lower noise; at $\sigma_{n}^{2}/\mathrm{Hz}=\,$$110\text{\,}\mathrm{dBm}$/Hz, training may even diverge.
+The corresponding accuracy loss due to clipping (considering the peak power constraint) is shown in Fig.
+6.
+At the later stage of training, the loss of test accuracy is about 1.5–2%, while around 50 rounds it can peak at up to 8% for the drop of accuracy mentioned previously.
 (a) Single-carrier.
@@
 This work highlights an overlooked practical issue in over-the-air computation-federated learning: the instantaneous peak power constraint imposed by the nonlinearity in power amplifiers.
-This is an issue in over-the-air computation-federated learning since it uses analog transmission and some works consider multi-carrier OFDM transmissions.
-We applied the classic PAPR reduction methods of iterative amplitude clipping and frequency-domain filtering.
-Enforcing peak power constraint can indeed degrade aggregation and final accuracy, in some cases, the effects are marginal, and in some other cases (multi-carrier with low noise level), cause divergence.
+This is a practical issue in over-the-air computation-federated learning since amplitude-modulated analog transmission is by itself a potentially high peak-to-average power ratio signal (depending on the distributions of values to be transmitted) and multi-carrier over-the-air computation-OFDM would inherit the high PAPR issue of OFDM.
+As the first work that reveals this problem in AirComp-FL, in order to confirm that this is actually a problem to be solved, we applied the most classic PAPR reduction method of iterative amplitude clipping and frequency-domain filtering that constrains the signal within its PAPR limits without causing distortions to adjacent frequency bands.
+We observe that enforcing peak power constraint can indeed degrade aggregation and final accuracy, in some cases, the effects are marginal, and in some other cases (multi-carrier with low noise level), cause divergence.
 Counterintuitively, the impact is stronger at lower noise for multi-carrier OFDM because the scaling factor from optimization is higher and therefore the in-band distortion is too much amplified.